# Talent Hoarding in Organizations

### v2

@@ 2.1 Evidence from an Employee Survey within the Firm
To provide the first detailed evidence on the dynamics of talent hoarding, I conduct a large-scale survey at the firm I study. All employees in my sample were invited via e-mail by the firm’s HR department and were asked to provide their perspectives on the internal labor market at the firm. Employees described challenges regarding their internal career progression both in the form of multiple-choice answers and in free-text responses. [^3] For questions requiring free-text responses, I exclude respondents from the analysis who did not provide any answer. An abbreviated version of the survey instrument is presented in Appendix Section D.2. The survey received over 15,000 responses, yielding a 50% response rate. Respondents are similar to the overall population in terms of their demographics (Appendix Table D1). [^4] Because of strict data protection regulations in Germany, I am not able to link the survey responses to the administrative employee records at the individual level. This limitation precludes a direct comparison of respondents to non-respondents in the administrative data. Respondents report a variety of different actions through which managers hoard talent, which include suppressing public signals of worker ability, restricting access to high-visibility projects or trainings, and explicitly discouraging workers from applying to internal positions. Appendix Table B1 provides selected quotations. In addition, when asked to state the biggest challenge to their internal career progression, the modal answer (provided by 22% of workers who respond to this question) is managers’ limited support for career progression, such as refusal to assist in career planning and restricted access to trainings that would increase workers’ visibility outside the team. Managers’ actions appear to strongly influence workers’ application decisions. 41% of respondents indicate that they are afraid to apply to internal positions, fearing negative repercussions if managers find out about the application. 25% of workers state that they feel the need to ask for managers’ permission before applying for an internal job opening, even though the firm’s internal policies are meant to enable workers to initiate an application on their own. 16% of respondents indicate that applying away from the team is seen as disloyal. These findings suggest that fear of retaliation represents a key dimension through which talent hoarding deters worker applications.
@@ 2.2 A Simple Framework of Talent Hoarding
To formally define talent hoarding, consider a firm that employs two types of agents, managers $m$ and workers $i$. For simplicity, teams are composed of one manager and one worker. Workers are characterized by a marginal productivity $\alpha_{i}$ drawn from some known distribution $G$. The firm seeks to efficiently allocate talent to maximize total firm productivity by choosing which workers to promote to managerial positions. Consistent with Rosen (1982), productivity is maximized when high-ability workers are assigned to high-level positions. Thus, in the absence of any constraints, the firm would fill a new managerial vacancy with the most productive worker and would fill the worker’s vacated position by hiring a worker from outside of the firm (i.e., a random draw from $G$). In practice, firms neither perfectly observe worker productivity, nor do they know which workers would accept a promotion. Accordingly, managers are tasked with identifying high-productivity workers and encouraging them to seek promotions. However, if a high-productivity worker is promoted and leaves their team, that team incurs a productivity loss. Managers, who observe the productivity of their workers, are compensated according to team productivity, creating a conflict of interest between the firm and managers. Talent hoarding is defined as the actions taken by managers that lower the likelihood that a worker applies for and receives a promotion. [^5] Respondents to the survey discussed in Section 2.1 report that managers can deter workers from applying by explicitly discouraging or threatening them, underrating worker ability, and restricting access to high-visibility projects or training. In theory, the firm could offer managers a promotion-contingent contract to resolve the misaligned incentives. In practice, firms generally do not compensate managers for promoting their workers, plausibly because of the practical challenges associated with these contracts (discussed in detail in Friebel and Raith, 2022). The one-period framework proceeds as follows. A managerial vacancy opens exogenously. $M$ managers observe the productivity of the worker on their teams and decide to what extent to engage in talent hoarding. Based on managers’ choices, workers decide whether to apply for a promotion. The firm observes noisy signals of worker productivity (e.g., by conducting interviews) and chooses the worker with the highest signal to fill the vacancy. The promoted worker’s previous position is replaced with an outside hire. Team productivity is realized and managers are compensated. Let $\beta\in[0,\infty)$ index the degree of talent hoarding chosen by a manager, with 0 representing no talent hoarding. [^6] Survey responses presented in Appendix Table B1 indicate that suppression of potential ratings and pressure to refrain from applying are common examples of talent hoarding that can be represented by $\beta$. In Section 4.2, I construct a direct measure of talent hoarding by comparing the private performance ratings to the public potential ratings that managers assign to workers, where $\beta$ can be interpreted as reflecting the disparity between these signals. Let $q(\alpha_{i},\beta)$ denote the equilibrium probability that a worker with productivity $\alpha_{i}$ gets promoted, conditional on applying for a promotion. This promotion probability increases with worker productivity ($\frac{\partial q}{\partial\alpha_{i}}>0$), but decreases in the level of talent hoarding ($\frac{\partial q}{\partial\beta}<0$), and reflects the noisy signal received by the firm. Thus, one can interpret talent hoarding as impacting workers through the likelihood that they get promoted. [^7] In practice, workers report that managers diminish their visibility, thus lowering their promotion prospects. In theory, talent hoarding can also operate through the cost of applying, which would yield similar predictions. Furthermore, I assume that the effect of talent hoarding on the conditional promotion probability is larger for more productive workers ($\frac{\partial^{2}q}{\partial\beta\partial\alpha_{i}}<0$). This assumption relates to the noisy signals of applicant productivity observed by the firm. Intuitively, since low-productivity workers are less likely than high-productivity workers to generate a favorable signal, there is less scope for talent hoarding to lower the likelihood that a low-productivity worker gets chosen among multiple applicants. One example in which this would be the case is if a firm employs a two-part screening strategy in which it chooses a subset of applicants to interview based on the CVs of all applicants. A very low-productivity worker may never clear the bar to be interviewed. Thus, the manager can do little to further lower their hiring likelihood. Workers decide whether to apply by weighing expected costs and benefits. Let $b$ denote the benefits of a successful application and $c$ denote the costs of applying. Workers apply if
@@
This definition yields following predictions for managers’ talent hoarding actions: Prediction 1. (Worker heterogeneity)
High-productivity workers (i.e., high $\alpha_{i}$) experience more talent hoarding. [^10] In many related models (e.g., Garicano, 2000), ability is multi-dimensional. To the extent that productivity on the worker’s current team $\alpha_{i}$ is positively related with performance external to the current team (e.g., leading a team), managers will have an incentive to hoard talent that conflicts with the firm’s objectives. Prediction 2. (Team heterogeneity)
Workers that are more difficult to replace (i.e., if $\bar{\alpha}$ is lower) experience more talent hoarding. Prediction 3. (Manager heterogeneity)
@@
In addition, workers’ decision rule implies that talent hoarding reduces the number of applicants and the quality of the applicant pool, limiting the firm’s ability to promote a high-productivity worker and generating misallocation of talent. Prediction 4. (Number of applicants)
@@
Talent hoarding has larger impacts for high-productivity workers. Therefore, higher levels of talent hoarding lead to a lower-quality applicant pool: If $\alpha_{1}<\alpha_{2}$ and $\beta_{1}<\beta_{2}$implies $\frac{\text{Pr}[i\text{ applies}|\alpha_{2},\beta_{1}]}{\text{Pr}[i\text{ applies}|\alpha_{1},\beta_{1}]}>\frac{\text{Pr}[i\text{ applies}|\alpha_{2},\beta_{2}]}{\text{Pr}[i\text{ applies}|\alpha_{1},\beta_{2}]}$
@@ 2.3 Implications for Empirical Tests
The predictions that follow from the conceptual framework suggest that the existence of talent hoarding can be inferred empirically from both manager and worker actions. First, the framework predicts that managers should engage more in talent hoarding if they face stronger incentives to hoard, for instance because of larger financial gains from hoarding. This prediction motivates empirical tests that relate managers’ revealed hoarding actions to variation in their hoarding incentives. This approach has the advantage that, relative to analyzing worker behavior, it is less likely to be confounded by potential worker inaction that may be driven by factors unrelated to talent hoarding. However, this test requires a direct measure of managerial talent hoarding actions, which is empirically challenging to construct given that talent hoarding involves hidden actions by definition. Moreover, since managerial hoarding can take many forms, any one measure can only capture a subset of talent hoarding actions. Second, the framework predicts that workers should be deterred from applying to positions outside the team if they are subject to talent hoarding. Accordingly, an empirical test based on workers requires both observing revealed worker actions and variation in workers’ exposure to managerial hoarding. Using high-stakes worker actions, such as internal applications or job transitions, [^11] Although the language used in the framework is centered around applications for promotions, the intuition and predictions also pertain to any type of internal application to a position outside of the team. From the manager’s perspective, both promotions and lateral moves have the same impacts on team productivity. From the firm’s perspective, deterring both types of transitions can create efficiency costs. has the advantage of being easier to observe in administrative data than managerial talent hoarding actions. In addition, such a test is not necessarily restricted to one specific measure of managerial hoarding behavior.
@@ 3.1 Firm Overview
I collect rich data on over 30,000 white-collar and management employees from a large manufacturing firm. This anonymous firm is one of the largest manufacturers in Europe and employs over 200,000 workers around the world, the plurality of which work in Germany. I restrict my sample to Germany because it represents the largest internal labor market at the firm. The firm operates in many other countries, including the United States. The firm’s establishments outside of Germany share many features in common with those in Germany, including organizational design and internal labor market policies (e.g., application systems, widespread use of performance and potential ratings). Because I am interested in career progression to higher-level positions, my analysis focuses on employees in white-collar and management positions (i.e., employees that are either already in or could ultimately be promoted to managerial positions). There are over 200 occupations represented in this sample, ranging from technical positions in engineering to support functions in HR and finance. Table 1 describes my analysis sample, which consists of over 300,000 employee-by-quarter observations from 2015 to 2018.121212To maintain confidentiality, I do not disclose the exact number of employees in my sample. Women represent 21% of employees in the sample, stemming from the underrepresentation of women in technical occupations. Employees’ educational qualifications are high, a result of restricting the sample to white-collar and management employees. The average employee holds a Bachelor’s degree and 92% of employees work full-time. Tenures at the firm are long, with an average of 13 years, highlighting the importance of internal career progression for employees’ long-term income. Managers (i.e., those that lead a team) comprise 19% of the sample. The demographics of the employees at the firm are comparable to other large manufacturing firms in Germany. In Appendix Table B3, I compare employees in my sample to those employees in large manufacturing firms in the BiBB, a representative survey of the German workforce conducted in 2018. I find very similar patterns with respect to most employee characteristics (e.g., gender, age, German citizenship, and marital status). The firm also resembles other large firms with respect to the design of its internal labor market (hkp, 2021). As in most large German firms, employees who want to switch to a new position or to be promoted are required to apply for the position using a centralized online job portal at the firm. All job openings are posted to the job portal, where openings are visible to all employees. Applications through the portal typically take less than five minutes to complete. While employees can choose to apply to multiple positions at the same time, the median applicant applies to only one position in a given quarter. Callback and hiring decisions are also recorded in the job portal. The internal labor market is comprised of competitive openings for new positions, much like the external labor market. Only a quarter of applications are successful and the internal labor market is both spatially and interpersonally diffuse. The firm operates in over 50 cities in 250 establishments throughout Germany and one-third of internal applications are for positions in a different city. For the vast majority of applications, applicants have not previously worked with the hiring manager of the position they are applying for, indicating that application and hiring decisions are distinct in the internal labor market. Because most teams are small, consisting of one manager and six workers on average, and because most workers in a team work in very similar positions, workers who want to move up the job ladder or seek different types of jobs must leave their team. At the firm, 97% of applications are to positions outside of a worker’s current team. Thus, managers who encourage their workers to develop their career lose team members. While workers can independently apply to any internal job opening, managers’ are tasked with assisting employees in their career planning (e.g., through regular conversations, mentoring activities, assignments to trainings). However, these activities are neither directly incentivized nor monitored at the firm. The performance of all employees, including managers themselves, is regularly evaluated by their direct supervisor. Individual contributors are only rated based on their individual achievements, but manager performance ratings take into account the performance of the team. While performance ratings have modest impacts on bonus payments for workers at lower levels, the performance of the manager and her team has substantial impacts on manager pay. For instance, for managers who hold executive positions (e.g., department head or above), bonus payments represent a quarter of their total compensation. Therefore, the composition and performance of the team can have important effects on manager compensation.
@@ 3.2 Personnel Records and Application Data
My data combine the firm’s internal personnel records from 1998 to 2019 with the universe of application and hiring decisions from 2015 to 2019. The personnel records capture over 30,000 employees, and the application data cover over 16,000 job openings and over 200,000 external and internal applicants. In my main analysis sample, I restrict to employees active at the firm from 2015 to 2018, for whom I can observe outcomes through 2019, yielding a sample of over 300,000 employee-by-quarter observations. Appendix Section D.1 contains details on data construction. The personnel records contain a large set of employee characteristics, such as age, educational qualifications, and tenure. The records also contain detailed position characteristics, such as position titles, leadership responsibility, and the reporting distance to the CEO. I supplement this data with payroll data, capturing employees’ working hours, earnings, and bonus payments. The richness of these data allow me to account for factors unrelated to talent hoarding that may influence workers’ career progression. Unless otherwise noted, all analyses include the following set of controls: female, age, German citizenship, educational qualifications, marital status, family status, parental leave, firm tenure, function, division, location, full-time status, hours, and number of direct reports. In addition, the personnel records capture workers’ assignments to teams and managers. I use this information to characterize manager behavior and construct measures of manager quality based on past outcomes of managers’ team members (e.g., absenteeism, turnover). Because these data capture team assignments over time, I am also able to construct measures of managers’ formal ties to other units at the firm, by measuring whether they have previously worked in that unit. Finally, I collect information from the firm’s talent management system that allow me to measure manager actions. These include worker evaluations, such as performance and potential ratings, which are conducted by a worker’s direct supervisor. The data also capture manager actions with respect to assigning workers to internal trainings that likely increase visibility outside the team. When testing how managerial talent hoarding affects worker actions, I focus on worker applications as a key outcome. Because all job openings are posted to the centralized portal and all applications are required to be submitted through the portal, I am able to construct a panel dataset of employees’ application histories at the firm from 2015 to 2018. Through the job portal, I also observe the outcome of each application in terms of rejections, interview callbacks, and subsequent hiring outcomes. Separately measuring applications and hiring outcomes is important given that workers’ forgone job switches could be explained by low hiring success unrelated to talent hoarding.

### v3

@@ 2.1 Evidence from an Employee Survey within the Firm
To provide the first detailed evidence on the dynamics of talent hoarding, I conduct a large-scale survey at the firm I study. Employees described challenges regarding their internal career progression both in the form of multiple-choice answers and in free-text responses. [^3] For questions requiring free-text responses, I exclude respondents from the analysis who did not provide any answer. The survey received over 15,000 responses, yielding a 50% response rate. Section 3.3 provides details about the survey implementation and response patterns. Appendix Section D.2 presents the survey instrument.
Respondents report a variety of different actions through which managers hoard talent, which include suppressing public signals of worker ability, restricting access to high-visibility projects or trainings, and explicitly discouraging workers from applying to internal positions. Appendix Table B1 provides selected quotations. In addition, when asked to state the biggest challenge to their internal career progression, the modal answer (provided by 22% of workers who respond to this question) is managers’ limited support for career progression, such as refusal to assist in career planning and restricted access to trainings that would increase workers’ visibility outside the team.
Managers’ actions appear to strongly influence workers’ application decisions. 41% of respondents indicate that they are afraid to apply to internal positions, fearing negative repercussions if managers find out about the application. 25% of workers state that they feel the need to ask for managers’ permission before applying for an internal job opening, even though the firm’s internal policies are meant to enable workers to initiate an application on their own. 16% of respondents indicate that applying away from the team is seen as disloyal. These findings suggest that fear of retaliation represents a key dimension through which talent hoarding deters worker applications.
@@ 2.2 A Simple Framework of Talent Hoarding
To formally define talent hoarding, consider a firm that employs two types of agents, managers $m$ and workers $i$. For simplicity, teams are composed of one manager and one worker. Workers are characterized by a marginal productivity $\alpha_{i}$ drawn from some known distribution $G$. The firm seeks to efficiently allocate talent to maximize total firm productivity by choosing which workers to promote to managerial positions. Consistent with Rosen (1982), productivity is maximized when high-ability workers are assigned to high-level positions. Thus, in the absence of any constraints, the firm would fill a new managerial vacancy with the most productive worker and would fill the worker’s vacated position by hiring a worker from outside of the firm (i.e., a random draw from $G$).
In practice, firms neither perfectly observe worker productivity, nor do they know which workers would accept a promotion. Accordingly, managers are tasked with identifying high-productivity workers and encouraging them to seek promotions. However, if a high-productivity worker is promoted and leaves their team, that team incurs a productivity loss. Managers, who observe the productivity of their workers, are compensated according to team productivity, creating a conflict of interest between the firm and managers. Talent hoarding is defined as the actions taken by managers that lower the likelihood that a worker applies for and receives a promotion. [^4] Respondents to the survey discussed in Section 2.1 report that managers can deter workers from applying by explicitly discouraging or threatening them, underrating worker ability, and restricting access to high-visibility projects or training. In theory, the firm could offer managers a promotion-contingent contract to resolve the misaligned incentives. In practice, firms generally do not compensate managers for promoting their workers, plausibly because of the practical challenges associated with these contracts (discussed in detail in Friebel and Raith, 2022).
The one-period framework proceeds as follows. A managerial vacancy opens exogenously. $M$ managers observe the productivity of the worker on their teams and decide to what extent to engage in talent hoarding. Based on managers’ choices, workers decide whether to apply for a promotion. The firm observes noisy signals of worker productivity (e.g., by conducting interviews) and chooses the worker with the highest signal to fill the vacancy. The promoted worker’s previous position is replaced with an outside hire. Team productivity is realized and managers are compensated.
@@
Let $\beta\in[0,\infty)$ index the degree of talent hoarding chosen by a manager, with 0 representing no talent hoarding. [^5] Survey responses presented in Appendix Table B1 indicate that suppression of potential ratings and pressure to refrain from applying are common examples of talent hoarding that can be represented by $\beta$. In Section 4.2, I construct a direct measure of talent hoarding by comparing the private performance ratings to the public potential ratings that managers assign to workers, where $\beta$ can be interpreted as reflecting the disparity between these signals. Let $q(\alpha_{i},\beta)$ denote the equilibrium probability that a worker with productivity $\alpha_{i}$ gets promoted, conditional on applying for a promotion. This promotion probability increases with worker productivity ($\frac{\partial q}{\partial\alpha_{i}}>0$), but decreases in the level of talent hoarding ($\frac{\partial q}{\partial\beta}<0$), and reflects the noisy signal received by the firm. Thus, one can interpret talent hoarding as impacting workers through the likelihood that they get promoted. [^6] In practice, workers report that managers diminish their visibility, thus lowering their promotion prospects. In theory, talent hoarding can also operate through the cost of applying, which would yield similar predictions. Furthermore, I assume that the effect of talent hoarding on the conditional promotion probability is larger for more productive workers ($\frac{\partial^{2}q}{\partial\beta\partial\alpha_{i}}<0$). This assumption relates to the noisy signals of applicant productivity observed by the firm. Intuitively, since low-productivity workers are less likely than high-productivity workers to generate a favorable signal, there is less scope for talent hoarding to lower the likelihood that a low-productivity worker gets chosen among multiple applicants. One example in which this would be the case is if a firm employs a two-part screening strategy in which it chooses a subset of applicants to interview based on the CVs of all applicants. A very low-productivity worker may never clear the bar to be interviewed. Thus, the manager can do little to further lower their hiring likelihood.
Workers decide whether to apply by weighing expected costs and benefits. Let $b$ denote the benefits of a successful application and $c$ denote the costs of applying. Workers apply if
@@
It follows that talent hoarding reduces the probability that workers leave the team (i.e., $\frac{\partial p}{\partial\beta}<0$). [^7] For simplicity, this framework does not distinguish between different worker types. Talent hoarding may exacerbate between-group differences in promotions if it has a differential effect on workers’ application decisions.
@@
This definition yields following predictions for managers’ talent hoarding actions:
Prediction 1. (Worker heterogeneity)
High-productivity workers (i.e., high $\alpha_{i}$) experience more talent hoarding. [^9] In many related models (e.g., Garicano, 2000), ability is multi-dimensional. To the extent that productivity on the worker’s current team $\alpha_{i}$ is positively related with performance external to the current team (e.g., leading a team), managers will have an incentive to hoard talent that conflicts with the firm’s objectives.
Prediction 2. (Team heterogeneity)
Workers that are more difficult to replace (i.e., if $\bar{\alpha}$ is lower) experience more talent hoarding.
Prediction 3. (Manager heterogeneity)
@@
In addition, workers’ decision rule implies that talent hoarding reduces the number of applicants and the quality of the applicant pool, limiting the firm’s ability to promote a high-productivity worker and generating misallocation of talent.
Prediction 4. (Number of applicants)
If workers face less talent hoarding, they are more likely to apply for a promotion.
@@
Talent hoarding has larger impacts for high-productivity workers. Therefore, higher levels of talent hoarding lead to a lower-quality applicant pool:
If $\alpha_{1}<\alpha_{2}$ and $\beta_{1}<\beta_{2}$implies $\frac{\text{Pr}[i\text{ applies}|\alpha_{2},\beta_{1}]}{\text{Pr}[i\text{ applies}|\alpha_{1},\beta_{1}]}>\frac{\text{Pr}[i\text{ applies}|\alpha_{2},\beta_{2}]}{\text{Pr}[i\text{ applies}|\alpha_{1},\beta_{2}]}$
@@ 2.3 Implications for Empirical Tests
The predictions that follow from the conceptual framework suggest that the existence of talent hoarding can be inferred empirically from both manager and worker actions.
First, the framework predicts that managers should engage more in talent hoarding if they face stronger incentives to hoard, for instance because of larger financial gains from hoarding. This prediction motivates empirical tests that relate managers’ revealed hoarding actions to variation in their hoarding incentives. This approach has the advantage that, relative to analyzing worker behavior, it is less likely to be confounded by potential worker inaction that may be driven by factors unrelated to talent hoarding. However, this test requires a direct measure of managerial talent hoarding actions, which is empirically challenging to construct given that talent hoarding involves hidden actions by definition. Moreover, since managerial hoarding can take many forms, any one measure can only capture a subset of talent hoarding actions.
Second, the framework predicts that workers should be deterred from applying to positions outside the team if they are subject to talent hoarding. Accordingly, an empirical test based on workers requires both observing revealed worker actions and variation in workers’ exposure to managerial hoarding. Using high-stakes worker actions, such as internal applications or job transitions, [^10] Although the language used in the framework is centered around applications for promotions, the intuition and predictions also pertain to any type of internal application to a position outside of the team. From the manager’s perspective, both promotions and lateral moves have the same impacts on team productivity. From the firm’s perspective, deterring both types of transitions can create efficiency costs. has the advantage of being easier to observe in administrative data than managerial talent hoarding actions. In addition, such a test is not necessarily restricted to one specific measure of managerial hoarding behavior.
@@ 3.1 Firm Overview
I collect rich data on over 30,000 white-collar and management employees from a large manufacturing firm. This anonymous firm is one of the largest manufacturers in Europe and employs over 200,000 workers around the world, the plurality of which work in Germany.
I restrict my sample to Germany because it represents the largest internal labor market at the firm. The firm operates in many other countries, including the United States. The firm’s establishments outside of Germany share many features in common with those in Germany, including organizational design and internal labor market policies (e.g., application systems, widespread use of performance and potential ratings). Because I am interested in career progression to higher-level positions, my analysis focuses on employees in white-collar and management positions (i.e., employees that are either already in or could ultimately be promoted to managerial positions). There are over 200 occupations represented in this sample, ranging from technical positions in engineering to support functions in HR and finance.
Table 1 describes my analysis sample, which consists of over 300,000 employee-by-quarter observations from 2015 to 2018.111111To maintain confidentiality, I do not disclose the exact number of employees in my sample. Women represent 21% of employees in the sample, stemming from the underrepresentation of women in technical occupations. Employees’ educational qualifications are high, a result of restricting the sample to white-collar and management employees. The average employee holds a Bachelor’s degree and 92% of employees work full-time. Tenures at the firm are long, with an average of 13 years, highlighting the importance of internal career progression for employees’ long-term income. Managers (i.e., those that lead a team) comprise 19% of the sample.
The demographics of the employees at the firm are comparable to other large manufacturing firms in Germany. In Appendix Table B3, I compare employees in my sample to those employees in large manufacturing firms in the BiBB, a representative survey of the German workforce conducted in 2018. I find very similar patterns with respect to most employee characteristics (e.g., gender, age, German citizenship, and marital status).
The firm also resembles other large firms with respect to the design of its internal labor market (hkp, 2021). As in most large German firms, employees who want to switch to a new position or to be promoted are required to apply for the position using a centralized online job portal at the firm. All job openings are posted to the job portal, where openings are visible to all employees. Applications through the portal typically take less than five minutes to complete. While employees can choose to apply to multiple positions at the same time, the median applicant applies to only one position in a given quarter. Callback and hiring decisions are also recorded in the job portal.
The internal labor market is comprised of competitive openings for new positions, much like the external labor market. Only a quarter of applications are successful and the internal labor market is both spatially and interpersonally diffuse. The firm operates in over 50 cities in 250 establishments throughout Germany and one-third of internal applications are for positions in a different city. For the vast majority of applications, applicants have not previously worked with the hiring manager of the position they are applying for, indicating that application and hiring decisions are distinct in the internal labor market.
Because most teams are small, consisting of one manager and six workers on average, and because most workers in a team work in very similar positions, workers who want to move up the job ladder or seek different types of jobs must leave their team. At the firm, 97% of applications are to positions outside of a worker’s current team. Thus, managers who encourage their workers to develop their career lose team members. While workers can independently apply to any internal job opening, managers’ are tasked with assisting employees in their career planning (e.g., through regular conversations, mentoring activities, assignments to trainings). However, these activities are neither directly incentivized nor monitored at the firm.
The performance of all employees, including managers themselves, is regularly evaluated by their direct supervisor. Individual contributors are only rated based on their individual achievements, but manager performance ratings take into account the performance of the team. While performance ratings have modest impacts on bonus payments for workers at lower levels, the performance of the manager and her team has substantial impacts on manager pay. For instance, for managers who hold executive positions (e.g., department head or above), bonus payments represent a quarter of their total compensation. Therefore, the composition and performance of the team can have important effects on manager compensation.
@@ 3.2 Personnel Records and Application Data
My data combine the firm’s internal personnel records from 1998 to 2019 with the universe of application and hiring decisions from 2015 to 2019. The personnel records capture over 30,000 employees, and the application data cover over 16,000 job openings and over 200,000 external and internal applicants. In my main analysis sample, I restrict to employees active at the firm from 2015 to 2018, for whom I can observe outcomes through 2019, yielding a sample of over 300,000 employee-by-quarter observations. Appendix Section D.1 contains details on data construction.
The personnel records contain a large set of employee characteristics, such as age, educational qualifications, and tenure. The records also contain detailed position characteristics, such as position titles, leadership responsibility, and the reporting distance to the CEO. I supplement this data with payroll data, capturing employees’ working hours, earnings, and bonus payments. The richness of these data allow me to account for factors unrelated to talent hoarding that may influence workers’ career progression. Unless otherwise noted, all analyses include the following set of controls: female, age, German citizenship, educational qualifications, marital status, family status, parental leave, firm tenure, function, division, location, full-time status, hours, and number of direct reports.
@@
In addition, the personnel records capture workers’ assignments to teams and managers. I use this information to characterize manager behavior and construct measures of manager quality based on past outcomes of managers’ team members (e.g., absenteeism, turnover). Because these data capture team assignments over time, I am also able to construct measures of managers’ formal ties to other units at the firm, by measuring whether they have previously worked in that unit. Finally, I collect information from the firm’s talent management system that allow me to measure manager actions. These include worker evaluations, such as performance and potential ratings, which are conducted by a worker’s direct supervisor. The data also capture manager actions with respect to assigning workers to internal trainings that likely increase visibility outside the team.
When testing how managerial talent hoarding affects worker actions, I focus on worker applications as a key outcome. Because all job openings are posted to the centralized portal and all applications are required to be submitted through the portal, I am able to construct a panel dataset of employees’ application histories at the firm from 2015 to 2018. Through the job portal, I also observe the outcome of each application in terms of rejections, interview callbacks, and subsequent hiring outcomes. Separately measuring applications and hiring outcomes is important given that workers’ forgone job switches could be explained by low hiring success unrelated to talent hoarding.
3.3 Survey of Employees
To capture worker perceptions of career progression at the firm, I conducted a large-scale survey with the employees in my sample. Employees described challenges regarding their internal career progression both in multiple-choice answers and in the form of free-text responses. The survey implementation followed the standard protocol that the firm uses for internal surveys. The survey was approved by the firm’s data protection officer and the workers’ council and was conducted through the firm’s online survey tool. Employees were invited via e-mail by the firm’s HR department. The invitation introduced the survey as eliciting employee feedback about the internal labor market to improve HR practices and provided employees with a participation link. See Appendix Section D.2 for information on more details on the implementation and the survey instrument.
The survey received over 15,000 responses, yielding a 50% response rate. Respondents are similar to the overall population in terms of their demographics (Appendix Table D1). [^12] Because of strict data protection regulations in Germany, I am not able to link the survey responses to the administrative employee records at the individual level. This limitation precludes a direct comparison of respondents to non-respondents in the administrative data. As is common for surveys within organizations, the survey did not include financial incentives. However, the survey invitation highlighted that employees’ responses would be used to improve the internal labor market. To increase trust in the survey, the survey invitation mentioned that the survey was approved by the workers’ council. Additionally, before the survey launch, all HR employees in Germany were briefed and instructed to address any employee inquiries, ensuring clarity and transparency about the survey’s purpose and approval.